OpenStudy (anonymous):

using substitution: y=3x+6 4x+5y=11

OpenStudy (lgbasallote):

use substitution...since y = 3x + 6 you can substitute that into the y of 4x + 5y = 11 it will look like this: \[4x + 5(3x + 6) = 11\] \[4x + 15x +30 = 11\] 19x = 11 - 30\] and so on

OpenStudy (lgbasallote):

just solve for x then when you have your x you can solve for y by substitution

OpenStudy (anonymous):

can you just give me the answer...PLEASE

OpenStudy (anonymous):

y=3 x=-1

OpenStudy (anonymous):

x=-1 y=3

myininaya (myininaya):

@ahaines14 and @hogr-eng Please don't just give the answers. It does not benefit anyone. OpenStudy's mission is to be an educational site not an answer providing site. We want to help people. It is against our CoC to just give answers. Try to engage the asker a bit please. Try to let them come to the answer. Try to teach. @brandonboo Please don't just ask for answers. This is also against our CoC. Please follow the CoC: http://openstudy.com/code-of-conduct